# Stein-Rule Shrinkage for Stochastic Gradient Estimation in High Dimensions

### v1

@@ Abstract
Stochastic gradient methods are central to large-scale learning, yet their analysis typically treats mini-batch gradients as unbiased estimators of the population gradient. In high-dimensional settings, however, classical results from statistical decision theory show that unbiased estimators are generally inadmissible under quadratic loss, suggesting that standard stochastic gradients may be suboptimal from a risk perspective. In this work, we formulate stochastic gradient computation as a high-dimensional estimation problem and introduce a decision-theoretic framework based on Stein-rule shrinkage. We construct a shrinkage gradient estimator that adaptively contracts noisy mini-batch gradients toward a stable restricted estimator derived from historical momentum. The shrinkage intensity is determined in a data-driven manner using an online estimate of gradient noise variance, leveraging second-moment statistics commonly maintained by adaptive optimization methods. Under a Gaussian noise model and for dimension $p\geq 3$, we show that the proposed estimator uniformly dominates the standard stochastic gradient under squared error loss and is minimax-optimal in the classical decision-theoretic sense. We further demonstrate how this estimator can be incorporated into the Adam optimizer, yielding a practical algorithm with negligible additional computational cost. Empirical evaluations on CIFAR10 and CIFAR100, across multiple levels of label noise, show consistent improvements over Adam in the large-batch regime. Ablation studies indicate that the gains arise primarily from selectively applying shrinkage to high-dimensional convolutional layers, while indiscriminate shrinkage across all parameters degrades performance. These results illustrate that classical shrinkage principles provide a principled and effective approach to improving stochastic gradient estimation in modern deep learning.
@@ 3 Theoretical Results
Refer to Appendix B for the proof.
@@
For the proof, refer to Appendix B.
@@
Refer to Appendix B for the proof.
@@
For the proof, refer to Appendix B.
@@ 5 Experimental Results
We present a controlled empirical study designed to isolate the optimizer’s effect while keeping the rest of the training pipeline fixed. Our experiments span the CIFAR10 and CIFAR100 datasets under three levels of label noise: 0%, 5%, and 10% (0.0, 0.05, 0.1). Unless otherwise stated, the optimizer is the only varying component; the backbone, preprocessing, number of epochs, batch size, and evaluation metrics are held constant to ensure a fair comparison. The complete implementation is available at: https://github.com/mamintoosi-papers-codes/SR-Adam.
@@ 5.1 Implementation Details of SR-Adam
  • The algorithm operates on a per-parameter-group basis, enabling heterogeneous behavior across different layers. For instance, Stein-rule shrinkage is selectively applied only to convolutional layer weights –where gradient estimates are highly dimensional and susceptible to noise– while standard Adam updates are used for fully connected layers and biases. This design choice is empirically validated in our ablation study (Section A.2) and aligns with the observed benefits of Stein correction in high-dimensional settings.
@@ 5.2 Model Architecture and SR-Adam Application
We employ a lightweight SimpleCNN backbone for all experiments to focus on the optimizer’s behavior without architectural confounds. SR-Adam selectively applies the Stein-rule shrinkage component only to the weights of the convolutional layers, while standard Adam updates are used for all other parameters, including fully-connected layers and biases. This strategy is motivated by the high-dimensional nature of convolutional gradients, where Stein-type estimators offer the most benefit by reducing noise, while the lower-dimensional fully-connected layers are better served by the direct adaptive updates of standard Adam. We empirically validate this design choice in our ablation study (Section A.2), where applying shrinkage to all layers did not yield additional performance gains.
@@ 5.4 Computational Cost
All experiments were conducted on a single desktop workstation equipped with an Intel Core i3-9100 CPU, 16 GB of system RAM, and an NVIDIA GeForce RTX 3090 GPU (24 GB VRAM). The implementation was in PyTorch. A complete multi-run grid over both CIFAR-10 and CIFAR-100 –5 optimizers (including one method mentioned in Appendix), 3 label-noise levels, batch size 512, 5 independent seeds, 20 epochs each– concluded in approximately 1,450 minutes (24 hours).
@@
5.5 Experimental Results
As mentioned above, we evaluate SR-Adam against some baseline methods, on two famous dataset using a Simple CNN backbone. To probe robustness to label noise, we add corruptions at three levels: 0.0, 0.05, and 0.1. For each dataset/noise combination, we conduct 5 independent runs with different random seeds and report mean $\pm$ std across runs. Accuracy is reported as a percentage (higher is better); loss is minimized (lower is better). In method comparison tables (3 and 4), we bold the best entry per dataset/noise column according to the respective metric direction.
@@ 5.5 Experimental Results
Figure 1: Bar chart comparison of best test accuracy on CIFAR10 under different label noise levels (0.0, 0.05, 0.1). Each subplot corresponds to one noise level. Bars show mean accuracy over five independent runs, and error bars indicate one standard deviation.
Figure 2: Bar chart comparison of best test accuracy on CIFAR100 under different label noise levels (0.0, 0.05, 0.1). Despite the increased task difficulty, SR-Adam consistently outperforms baseline optimizers, particularly in noisy regimes.
These visualizations complement Table 3 by highlighting consistent performance trends across noise regimes. In particular, SR-Adam achieves the highest mean accuracy in all noise settings for both datasets, with the performance gap becoming more pronounced as label noise increases. This trend supports the hypothesis that Stein-rule shrinkage is especially effective in stabilizing noisy gradient estimates under corrupted supervision.
@@
All comparisons use an identical training protocol and data processing across methods; only the optimizer changes. We hold constant the backbone, number of epochs, batch size, label-noise injection, evaluation procedure, and seed schedule (five seeds per configuration), while using standard, fixed hyperparameters per optimizer throughout. To avoid discrepancies due to third-party implementations, all optimizers are implemented within a single, consistent codebase with matching interfaces, and the full executable code is publicly available from the paper repository  [^1] https://github.com/mamintoosi-papers-codes/SR-Adam.
@@ 5.5.3 Qualitative Analysis: Prediction Visualization
Figure 8: Qualitative comparison on CIFAR100 with 5% label noise. Similar to Figure 7, this visualization shows sample predictions from the best-performing checkpoints. CIFAR100’s 100-class taxonomy presents a significantly harder recognition task than CIFAR10. The fine-grained categories (e.g., distinguishing between different tree species or vehicle types) require more nuanced feature learning. SR-Adam’s shrinkage mechanism helps stabilize gradient estimates in this high-noise, high-complexity regime, leading to more robust decision boundaries. The confidence scores reveal that SR-Adam produces higher-confidence predictions on correctly classified samples, suggesting improved calibration and reduced uncertainty.
@@ 5.6 Statistical Significance Analysis
Table 5 summarises paired-sample $t$-tests (n=5n\!=\!5 runs) comparing the best test accuracy achieved by Adam and SR-Adam with a batch size of 512. On CIFAR-10, SR-Adam yields a consistent gain of 1.5–2.2 percentage points across all label-noise levels; the improvement is statistically significant (p<0.01p\!<\!0.01) in every case. On CIFAR-100 the margin is smaller (around 1 point), yet still significant for clean data and 10 % noise; the difference observed at 5 % noise does not reach significance (p=0.074p\!=\!0.074).
@@
6 Discussion and Conclusion
We further demonstrate how the proposed estimator can be instantiated within the Adam optimizer, yielding the SR-Adam algorithm. Empirical results indicate that the benefits of Stein-rule shrinkage are strongly regime-dependent. In particular, SR-Adam consistently improves upon Adam in the large-batch setting, where gradient noise is sufficiently moderate for shrinkage to be effective. These gains are especially pronounced under label noise, where stochastic gradients become increasingly unreliable.
Appendix A Ablation Study
@@
A.1 Effect of Batch Size
@@
A.2 Ablation Study on Shrinkage Scope
@@
The results, summarized in Table 7 and visualized in Figure 10, demonstrate a clear trend. Across all datasets and noise conditions, SR-Adam-All-Weights consistently underperforms both the Adam baseline and our selectively-applied SR-Adam. Notably, this performance degradation is present even in the noise-free setting, indicating that the issue is intrinsic to the unrestricted application of shrinkage rather than being a side effect of label noise.
@@
Appendix B Proof of the Main Results

### v2

@@ Abstract
Stochastic gradient methods are central to large-scale learning, but they treat mini-batch gradients as unbiased estimators, which classical decision theory shows are inadmissible in high dimensions. We formulate gradient computation as a high-dimensional estimation problem and introduce a framework based on Stein-rule shrinkage. We construct a gradient estimator that adaptively contracts noisy mini-batch gradients toward a stable estimator derived from historical momentum. The shrinkage intensity is determined in a data-driven manner using an online estimate of gradient noise variance, leveraging statistics from adaptive optimizers. Under a Gaussian noise model, we show our estimator uniformly dominates the standard stochastic gradient under squared error loss and is minimax-optimal. We incorporate this into the Adam optimizer, yielding SR-Adam, a practical algorithm with negligible computational cost. Empirical evaluations on CIFAR10 and CIFAR100 across multiple levels of input noise show consistent improvements over Adam in the large-batch regime. Ablation studies indicate that gains arise primarily from selectively applying shrinkage to high-dimensional convolutional layers, while indiscriminate shrinkage across all parameters degrades performance. These results illustrate that classical shrinkage principles provide a principled approach to improving stochastic gradient estimation in deep learning.
@@ 3 Theoretical Results
Refer to Appendix A for the proof.
@@
For the proof, refer to Appendix A.
@@
Refer to Appendix A for the proof.
@@
For the proof, refer to Appendix A.
@@ 5 Experimental Results
We present a controlled empirical study designed to isolate the optimizer’s effect while keeping the rest of the training pipeline fixed. Our experiments span the CIFAR10 and CIFAR100 datasets under three levels of input noise, corresponding to Gaussian standard deviations of 0.0, 0.05, and 0.1. Unless otherwise stated, the optimizer is the only varying component; the backbone, preprocessing, number of epochs, batch size, and evaluation metrics are held constant to ensure a fair comparison. The complete implementation is available at: https://github.com/mamintoosi-papers-codes/SR-Adam.
@@ 5.1 Implementation Details of SR-Adam
  • The algorithm operates on a per-parameter-group basis, enabling heterogeneous behavior across different layers. For instance, Stein-rule shrinkage is selectively applied only to convolutional layer weights –where gradient estimates are highly dimensional and susceptible to noise– while standard Adam updates are used for fully connected layers and biases. This design choice is empirically validated in our ablation study (Section 6.2) and aligns with the observed benefits of Stein correction in high-dimensional settings.
@@ 5.2 Model Architecture and SR-Adam Application
We employ a lightweight SimpleCNN backbone for all experiments to focus on the optimizer’s behavior without architectural confounds. SR-Adam selectively applies the Stein-rule shrinkage component only to the weights of the convolutional layers, while standard Adam updates are used for all other parameters, including fully-connected layers and biases. This strategy is motivated by the high-dimensional nature of convolutional gradients, where Stein-type estimators offer the most benefit by reducing noise, while the lower-dimensional fully-connected layers are better served by the direct adaptive updates of standard Adam. We empirically validate this design choice in our ablation study (Section 6.2), where applying shrinkage to all layers did not yield additional performance gains.
@@ 5.4 Computational Cost
All experiments were conducted on a single desktop workstation equipped with an Intel Core i3-9100 CPU, 16 GB of system RAM, and an NVIDIA GeForce RTX 3090 GPU (24 GB VRAM). The implementation was in PyTorch. A complete multi-run grid over both CIFAR-10 and CIFAR-100 –5 optimizers (including one method mentioned in ablation study), 3 input noise levels, batch size 512, 5 independent seeds, 20 epochs each– concluded in approximately 1,450 minutes (24 hours).
@@
5.5 Results and Analysis
As mentioned above, we evaluate SR-Adam against some baseline methods, on two famous dataset using a Simple CNN backbone. To probe robustness to input noise, we add Gaussian noise at three standard deviation levels: 0.0, 0.05, and 0.1. For each dataset/noise combination, we conduct 5 independent runs with different random seeds and report mean $\pm$ std across runs. Accuracy is reported as a percentage (higher is better); loss is minimized (lower is better). In method comparison tables (3 and 4), we bold the best entry per dataset/noise column according to the respective metric direction.
@@ 5.5 Results and Analysis
Figure 1: Bar chart comparison of best test accuracy on CIFAR10 under different input noise levels (std: 0.0, 0.05, 0.1). Each subplot corresponds to one noise level. Bars show mean accuracy over five independent runs, and error bars indicate one standard deviation.
Figure 2: Bar chart comparison of best test accuracy on CIFAR100 under different input noise levels (std: 0.0, 0.05, 0.1). Despite the increased task difficulty, SR-Adam consistently outperforms baseline optimizers, particularly in noisy regimes.
These visualizations complement Table 3 by highlighting consistent performance trends across noise regimes. In particular, SR-Adam achieves the highest mean accuracy in all noise settings for both datasets, with the performance gap becoming more pronounced as the level of input noise increases. This trend supports the hypothesis that Stein-rule shrinkage is especially effective in stabilizing noisy gradient estimates under corrupted supervision.
@@
All comparisons use an identical training protocol and data processing across methods; only the optimizer changes. We hold constant the backbone, number of epochs, batch size, input noise injection, evaluation procedure, and seed schedule (five seeds per configuration), while using standard, fixed hyperparameters per optimizer throughout. To avoid discrepancies due to third-party implementations, all optimizers are implemented within a single, consistent codebase with matching interfaces, and the full executable code is publicly available from the paper repository [^1] https://github.com/mamintoosi-papers-codes/SR-Adam.
@@ 5.5.3 Qualitative Analysis: Prediction Visualization
Figure 8: Qualitative comparison on CIFAR100 with an input noise standard deviation of 0.05. Similar to Figure 7, this visualization shows sample predictions from the best-performing checkpoints. CIFAR100’s 100-class taxonomy presents a significantly harder recognition task than CIFAR10. The fine-grained categories (e.g., distinguishing between different tree species or vehicle types) require more nuanced feature learning. SR-Adam’s shrinkage mechanism helps stabilize gradient estimates in this high-noise, high-complexity regime, leading to more robust decision boundaries. The confidence scores reveal that SR-Adam produces higher-confidence predictions on correctly classified samples, suggesting improved calibration and reduced uncertainty.
@@ 5.6 Statistical Significance Analysis
Table 5 summarises paired-sample $t$-tests (n=5n\!=\!5 runs) comparing the best test accuracy achieved by Adam and SR-Adam with a batch size of 512. On CIFAR-10, SR-Adam yields a consistent gain of 1.5–2.2 percentage points across all input noise levels; the improvement is statistically significant (p<0.01p\!<\!0.01) in every case. On CIFAR-100 the margin is smaller (around 1 point), yet still significant for clean data and 10 % noise; the difference observed at 5 % noise does not reach significance (p=0.074p\!=\!0.074).
@@
6 Ablation Study
@@
6.1 Effect of Batch Size
@@
6.2 Ablation Study on Shrinkage Scope
@@
The results, summarized in Table 7 and visualized in Figure 10, demonstrate a clear trend. Across all datasets and noise conditions, SR-Adam-All-Weights consistently underperforms both the Adam baseline and our selectively-applied SR-Adam. Notably, this performance degradation is present even in the noise-free setting, indicating that the issue is intrinsic to the unrestricted application of shrinkage rather than being a side effect of input noise.
@@
7 Discussion and Conclusion
We further demonstrate how the proposed estimator can be instantiated within the Adam optimizer, yielding the SR-Adam algorithm. Empirical results indicate that the benefits of Stein-rule shrinkage are strongly regime-dependent. In particular, SR-Adam consistently improves upon Adam in the large-batch setting, where gradient noise is sufficiently moderate for shrinkage to be effective. These gains are especially pronounced under input noise, where stochastic gradients become increasingly unreliable.
Appendix A Proof of the Main Results